OpenStudy (gabylovesyou):

Choose the slope-intercept equation of the line that passes through the point shown and is parallel to the line shown.

OpenStudy (gabylovesyou):

y = 1/3x + 14/3 y = 3x + 10 y = -3x - 2 y = - 1/3x + 10/3

OpenStudy (anonymous):

the point shown is -2,4

OpenStudy (anonymous):

can you tell the slope of the line ?

OpenStudy (gabylovesyou):

2, -8

OpenStudy (anonymous):

what 2,-8 ?

OpenStudy (gabylovesyou):

the slope

OpenStudy (anonymous):

slope is one value .. there is not coordinates for slope

OpenStudy (gabylovesyou):

then idk

OpenStudy (anonymous):

two options : find it using two points on the graph or graphic way : how much y changes if you go 1 unit towards the positive x direction

OpenStudy (gabylovesyou):

3

OpenStudy (anonymous):

two points on the graph are : (0,6) , (2,0) in order to find the slope you do : m = (y2- y1)/(x2-x1) = (0-6)/(2-0) = -3

OpenStudy (anonymous):

as you can see from the graph for every unit of x you go down by 3 units of y so graphic way gives us as well -3.

OpenStudy (anonymous):

now in order to get a line parallel to this line we have to have the same slope so now we have to construct a line equation from the given point -2,4 and a slope of -3 you do so by : y= y1 +m(x-x1)

OpenStudy (gabylovesyou):

y = 4 + -3(-2 -X1)

OpenStudy (anonymous):

you have to plug -2 in x1 and not in x

OpenStudy (anonymous):

just like you plugged 4 in y1

OpenStudy (gabylovesyou):

so it would be?

OpenStudy (anonymous):

y = 4 -3(x+2) y = 4 -3x -6 y = -3x -2

OpenStudy (gabylovesyou):

so its C?

OpenStudy (anonymous):

yes well in fact we could say it's C once we knew that the slope is -3

OpenStudy (gabylovesyou):

ok :)

OpenStudy (anonymous):

its the only option with such slope

OpenStudy (gabylovesyou):

thanks